Originally the cosumer face the budget line p2+x2 then tha price of good 1 double,the price of good 2 becomes 8 times larger and income become 4 time larger.write down an equation for the new budget line in terms of the original prices and income.


1
Expert's answer
2021-02-23T13:32:09-0500
Budget line 
Income =summation of (product price multiplied by the quantity )
M=p2+x2 
The new budget line will be
4M= 2p2+8x2 
M=0.5p2+2x2
The slope = 0.5/2 or 2/0.5
intercepts = M/2 and M/0.5